# Minimal length: a source of non-Hermiticity and non-locality in quantum mechanics

### v1

@@ Abstract
First, the study tries to shed light on the relationship between purely quantum mechanical momentum measurements (canonical momentum space) and measurements of the generalized momentum operator, including minimal length effects. Additionally, the existence of complex numbers in quantum mechanics seems justifiable as a consequence of minimal length. Finally, a novel method for generating quantum entangled states with complex quantum numbers inspired by the minimal length is also reported. Therefore, theories including a minimal length, like some quantum scenarios of gravity, seem to be able to enrich the current understanding of non-locality.
@@ I Introduction
Based on the Heisenberg Uncertainty Principle (HUP), there is no limit on the minimum value of length, provided that the information of momentum is completely lost. On the other hand, the role of gravity at quantum scales is mysterious Hossenfelder:2012jw ; Das:2008kaa ; Ali:2009zq ; Ali:2011fa . Nevertheless, considering Newtonian gravity in revisiting the Heisenberg microscope unveils a non-zero minimal length Hossenfelder:2012jw ; Mead:1964zz ; Mead:1966zz . It leads to the Generalized Uncertainty Principle (GUP), which is also reinforced in quantum scenarios of gravity Hossenfelder:2012jw ; Kempf:1994su . HUP lies at the heart of classically amazing phenomena like quantum non-locality Einstein:1935rr and correspondingly, the existence of minimal length and quantum aspects of gravity affect the quantum entanglement, the violation of Bell’s inequality Bose:2017nin ; Marletto:2017kzi ; Blado:2017gnf ; Aghababaei:2021yzx ; Aghababaei:2022rqi ; Park:2022jrs ; Moradpour:2023gsn ; Filho:2023vpx ; Marletto:2025fpm , and the information bounds Rastegin:2016xeb ; Pedram:2016gps ; Aghababaei:2022jxd . Indeed, we are even witnessing a fascinating claim that classical gravity produces quantum entanglement Aziz:2025ypo ; Marletto:2025fpm . Additionally, the conjectured correspondence between metric-affine gravity and quantum gravity scenarios Ali:2024tbd further motivates us to study the implications of the minimum length on various phenomena such as quantum non-locality and related topics.
@@ II Canonical momentum, generalized momentum, and complex numbers
In the quantum regime, we basically have $\hat{x},\hat{p}]=i\hbar$ and
@@
In the presence of minimal length, the canonical momentum operator $\hat{p}$ is replaced with the generalized momentum operator $\hat{P}(\equiv\hat{p}~(1+\beta\hat{p}^{2}))$ and the position operator remains unchanged (position representation Brau:1999uv ). Thus, we have $\hat{x},\hat{P}]=i\hbar(1+3\beta\hat{P}^{2})$ leading to Hossenfelder:2012jw
@@
that admits $\Delta x_{min}=\sqrt{3\beta}~\hbar\equiv\sqrt{\beta_{0}}~l_{p}$ where $l_{p}$ denotes the Planck length, respectively Hossenfelder:2012jw . In this regime, we have $\hat{H}=\frac{\hat{P}^{2}}{2m}$ for the Hamiltonian. Now, let us represent the eigenstate of $\hat{P}$ with $|P\rangle$ and thus $\hat{P}~|P_{0}\rangle=P_{0}~|P_{0}\rangle$. Moreover, since $\hat{P}$ and $\hat{p}$ commute, their info can be represented simultaneously. But, it should be noted that since $\hat{P}\neq\hat{p}$, their states are not the same ($|p\rangle\neq|P\rangle$), a point easily verifiable using $\langle x|P\rangle\neq\langle x|p\rangle$. Indeed,
@@
with these possibilities of $\mathcal{P}$
@@
$\mathcal{P}_{2}=-\frac{\mathcal{P}_{1}}{2}+i\frac{\sqrt{3}}{2}\bigg[\frac{\mathcal{P}_{1}}{2}+\frac{4}{\mathcal{A}}\bigg]\equiv e^{i\gamma},$
$\mathcal{P}_{3}=-\frac{\mathcal{P}_{1}}{2}-i\frac{\sqrt{3}}{2}\bigg[\frac{\mathcal{P}_{1}}{2}+\frac{4}{\mathcal{A}}\bigg]=e^{-i\gamma},$
$\gamma=-\arctan(\frac{\sqrt{3}}{2}+\frac{4\sqrt{3}}{\mathcal{A}\mathcal{P}_{1}}),$
@@
Interestingly enough, it is seen that Eq. (4) is also the eigenstates of the canonical momentum operator ($\hat{p}$) with eigenvalues $\mathcal{P}_{i}$. The latter means there is a threefold degeneracy for each value of $P$ that can be unveiled using $\hat{p}$. As the final remark, it is worth noting that the complex values of momentum and, indeed, the existence and implications of complex numbers in QM are a longstanding topic AMJP ; Renou:2021dvp ; Li:2021uof ; Chen:2021ril ; Albert:2023kah ; Lancaster:2022vgi . Therefore, complex values like $\mathcal{P}_{2}$ may convince us to conclude that the necessity of complex momentum (an old and still significant issue Albert:2023kah ) in quantum mechanics is a consequence of minimal length.
$|P\rangle$ and $|p\rangle$ denote the basis of the spaces of $\hat{P}$ and $\hat{p}$, respectively, and by considering the above argument, one can write the general solution of Eq. (3) as
@@
$\langle P_{1}|P_{2}\rangle=\sum_{l,m}\alpha_{l}^{\ast}\alpha_{m}\langle p_{l}|p_{m}\rangle\Rightarrow$ (9)
$\int\langle P_{1}|P_{2}\rangle\frac{dP_{2}}{1+\beta P_{2}^{2}}=\sum_{l,m}\int\alpha_{l}^{\ast}\alpha_{m}\langle p_{l}|p_{m}\rangle dp_{m},$
@@
$\sum_{l=1}^{3}|\alpha_{l}|^{2}=1.$ (10)
At this step, no constraint would make a difference between $\alpha_{l}$, and hence, the same factor is assumed for them all ($\alpha_{l}=\frac{1}{\sqrt{3}}$) from now on. Clearly, if measurement is done in the $|P\rangle$ bases, then we deal with only one state and thus one outcome. On the other hand, if a purely quantum-mechanical apparatus is employed (or equivalently, the measurement is performed in the $|p\rangle$ space), then three outcomes $\mathcal{P}_{i}$ are likely to occur after the collapse of the wave function.
@@ III Quantum entanglement production
$\mathcal{P}_{2}+\mathcal{P}_{3}=-\mathcal{P}_{1}\equiv\mathcal{P}_{1}^{\prime},$
$\mathcal{P}_{1}+\mathcal{P}_{3}=-\mathcal{P}_{2}\equiv\mathcal{P}_{2}^{\prime},$
$\mathcal{P}_{1}+\mathcal{P}_{2}=-\mathcal{P}_{3}\equiv\mathcal{P}_{3}^{\prime}.$ (11)
Comparing these results with Eq. (5) for which
@@
$\mathcal{P}_{j}^{\prime}\big[1+\beta(\mathcal{P}_{j}^{\prime})^{2}\big]=-\mathcal{P}_{j}(1+\beta\mathcal{P}_{j}^{2})=-P,$ (13)
@@
$|-P\rangle=\frac{1}{\sqrt{3}}\bigg(|\mathcal{P}_{1}^{\prime}\rangle+|\mathcal{P}_{2}^{\prime}\rangle+|\mathcal{P}_{3}^{\prime}\rangle\bigg),$
$=\frac{1}{\sqrt{3}}\bigg(|-\mathcal{P}_{1}\rangle+|-\mathcal{P}_{2}\rangle+|-\mathcal{P}_{3}\rangle\bigg).$ (14)
@@
At this step, consider a two-particle system whose total momentum is zero, so that $P_{1}=-P_{2}=P,$ leading to the momentum state $|P,-P\rangle$. Clearly, the outcome of measuring $P_{i}$ is completely predictable. The momentum conservation ($\sum_{j=1}^{j=2}P_{j}=0$) in the $\hat{P}-$space inspires the canonical momentum conservation ($\sum_{j=1}^{j=2}p_{j}=$0) generating the $p_{1}=-p_{2}$ condition. Interestingly enough, if one focuses on purely quantum mechanical momentum measurements ($|p\rangle$ space), then the state
$\frac{1}{\sqrt{3}}\bigg(|\mathcal{P}_{1},-\mathcal{P}_{1}\rangle+|\mathcal{P}_{2},-\mathcal{P}_{2}\rangle+|\mathcal{P}_{3},-\mathcal{P}_{3}\rangle\bigg),$ (15)
is dealt with, unveiling a novel quantum non-locality generation in the canonical momentum space. Therefore, the existence of minimal length may generate quantum non-locality. In this line, for a state like $\frac{1}{\sqrt{2}}\big[|P,-P\rangle+|-P,P\rangle\big]$, one reaches at
$\frac{1}{\sqrt{6}}\bigg(|\mathcal{P}_{1},-\mathcal{P}_{1}\rangle+|\mathcal{P}_{2},-\mathcal{P}_{2}\rangle+|\mathcal{P}_{3},-\mathcal{P}_{3}\rangle$ (16)
$+|-\mathcal{P}_{1},\mathcal{P}_{1}\rangle+|-\mathcal{P}_{2},\mathcal{P}_{2}\rangle+|-\mathcal{P}_{3},\mathcal{P}_{3}\rangle\bigg).$
As previously discussed, the same value for the coefficients has been assumed due to the lack of any other constraints on $\alpha_{k}$, thereby preserving simplicity.
@@ IV Summary and Concluding Remarks
The discovered degeneracy in $\langle x|P\rangle$ seems breakable using the canonical momentum operator. It also signals to us that minimal length (and thus GUP) may be an origin for the emergence of complex numbers in QM, which play impressive roles in various phenomena such as quantum non-locality AMJP ; Renou:2021dvp ; Li:2021uof ; Chen:2021ril ; Albert:2023kah ; Lancaster:2022vgi . Finally, and most importantly, calculations reveal the ability of a minimal length to create quantum entanglement. The simplest spin version of the EPR thought experiment Bohm:1957zz includes a two-$1/2$ spin particle system whose spin state in the total spin space is $|S=1,S_{z}=0\rangle$. The state’s decomposition using the spin state of each particle leads to the $\frac{1}{\sqrt{2}}\big[|+,-\rangle+|-,+\rangle\big]$ state addressing the quantum entanglement. The same behaviour is also obtained by writing $|P,-P\rangle$ in the canonical momentum bases of each particle, leading to Eq. (15). It is a direct signature of quantum non-locality generation due to the existence of a minimal length, so that the system state includes complex numbers.
Data availability Statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Declarations Conflict of interest
The authors declare no conflict of interest.

### v2

# Minimal length: A source of quantum non-locality

@@ Abstract
The narrow and subtle difference between the Hilbert spaces of operators corresponding to the purely quantum mechanical momentum and the generalized momentum that includes minimal length effects is polished. Additionally, the existence of complex numbers in quantum mechanics may be justifiable as a consequence of minimal length. A novel quantum entanglement generation is also reported indicating the power of theories including a minimal length in enriching the current understanding of quantum non-locality.
@@ I Introduction
Based on the Heisenberg Uncertainty Principle (HUP), there is no limit on the minimum value of length, provided that the momentum information is completely lost. On the other hand, the role of gravity at quantum scale is mysterious Hossenfelder:2012jw ; Das:2008kaa ; Ali:2009zq ; Ali:2011fa . Nevertheless, considering Newtonian gravity in revisiting the Heisenberg microscope unveils a non-zero minimal length Hossenfelder:2012jw ; Mead:1964zz ; Mead:1966zz . It leads to the Generalized Uncertainty Principle (GUP), which is also reinforced in quantum scenarios of gravity Hossenfelder:2012jw ; Kempf:1994su . HUP lies at the heart of classically amazing phenomena like quantum non-locality Einstein:1935rr and correspondingly, the existence of minimal length and quantum aspects of gravity affect the quantum entanglement, the violation of Bell’s inequality Bose:2017nin ; Marletto:2017kzi ; Blado:2017gnf ; Aghababaei:2021yzx ; Aghababaei:2022rqi ; Park:2022jrs ; Moradpour:2023gsn ; Filho:2023vpx ; Marletto:2025fpm , and the information bounds Rastegin:2016xeb ; Pedram:2016gps ; Aghababaei:2022jxd . Indeed, we are even witnessing a fascinating claim that classical gravity produces quantum entanglement Aziz:2025ypo ; Marletto:2025fpm . Additionally, the conjectured correspondence between metric-affine gravity and quantum gravity scenarios Ali:2024tbd further motivates us to study the implications of the minimum length on various phenomena such as quantum non-locality and related topics.
@@ II Canonical momentum, generalized momentum, and complex numbers
In the purely quantum mechanical regime, we basically have $\hat{x},\hat{p}]=i\hbar$ and
@@
In the presence of minimal length, the generalized momentum operator $\hat{P}(\equiv\hat{p}~(1+\beta\hat{p}^{2}))$ should be employed to measure momentum. It is also worthy to remember that, based on quantum mechanics, the canonical momentum operator ($\hat{p}$) is still a physical observable with its Hilbert space ($|p\rangle$). In the minimal length regime, the position operator remains unchanged (position representation Brau:1999uv ). We have $\hat{x},\hat{P}]=i\hbar(1+3\beta\hat{P}^{2})$ leading to Hossenfelder:2012jw
@@
that admits $\Delta x_{min}=\sqrt{3\beta}~\hbar\equiv\sqrt{\beta_{0}}~l_{p}$ where $l_{p}$ denotes the Planck length, respectively Hossenfelder:2012jw . In this regime, we have $\hat{H}=\frac{\hat{P}^{2}}{2m}$ for the Hamiltonian. Now, let us represent the eigenstate of $\hat{P}$ with $|P\rangle$ and thus $\hat{P}~|P_{0}\rangle=P_{0}~|P_{0}\rangle$. Moreover, since $\hat{P}$ and $\hat{p}$ commute, their info can be represented simultaneously. But, it should be noted that since $\hat{P}\neq\hat{p}$, their states (Hilbert spaces) are not the same ($|p\rangle\neq|P\rangle$), a point easily verifiable using $\langle x|P\rangle\neq\langle x|p\rangle$. Indeed,
@@
with these values of $\mathcal{P}$
@@
$\mathcal{P}_{2}=-\frac{\mathcal{P}_{1}}{2}+i\frac{\sqrt{3}}{2}\bigg[\frac{\mathcal{P}_{1}}{2}+\frac{4}{\mathcal{A}}\bigg],$
$\mathcal{P}_{3}=-\frac{\mathcal{P}_{1}}{2}-i\frac{\sqrt{3}}{2}\bigg[\frac{\mathcal{P}_{1}}{2}+\frac{4}{\mathcal{A}}\bigg],$
@@
Interestingly enough, it is seen that Eq. (4) is also the eigenstates of the canonical momentum operator ($\hat{p}$) with eigenvalues $\mathcal{P}_{i}$. Since $\beta$ is positive Hossenfelder:2012jw ; Kempf:1994su , when $\beta\rightarrow 0^{+}$, we have $\mathcal{P}_{1}\rightarrow P(\rightarrow p)$, $\mathcal{P}_{2}\rightarrow i\infty$, and $\mathcal{P}_{3}\rightarrow-i\infty$. Therefore, in the presence of minimal length, $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ have also real part ($\frac{\mathcal{P}_{1}}{2}$). In fact, whenever $\beta\neq 0$, while for each eigenvalue $p$ there is only one eigenvalue $P=p(1+\beta p^{2})$, the inverse is not necessarily always true as $p(1+\beta p^{2})=P$ is a third order equation of $p$. Hence, there is a threefold degeneracy for each value of $P$ that can be unveiled using $\hat{p}$. It is worth noting that the existence of complex values of momentum and indeed, the implications of complex numbers in QM are longstanding topics AMJP ; Renou:2021dvp ; Li:2021uof ; Chen:2021ril ; Albert:2023kah ; Lancaster:2022vgi . Consequently, the existence of complex momentum in quantum mechanics may be a consequence of minimal length.
$|P\rangle$ and $|p\rangle$ denote the basis of the Hilbert spaces corresponding to the operators $\hat{P}$ and $\hat{p}$, respectively, and by considering the above argument, one can write the general solution of Eq. (3) as
@@
$\langle P_{1}|P_{2}\rangle=\sum_{k,m}\alpha_{k}^{\ast}\alpha_{m}\langle p_{k}|p_{m}\rangle\Rightarrow$ (9)
$\int\langle P_{1}|P_{2}\rangle\frac{dP_{2}}{1+\beta P_{2}^{2}}=\sum_{k,m}\int\alpha_{k}^{\ast}\alpha_{m}\langle p_{k}|p_{m}\rangle dp_{m},$
@@
$\sum_{k=1}^{3}|\alpha_{k}|^{2}=1.$ (10)
Clearly, if measurement is done in the $|P\rangle$ bases, then we deal with only one state and thus one outcome. On the other hand, if a purely quantum-mechanical apparatus is employed (or equivalently, the measurement is performed in the $|p\rangle$ space), then three outcomes $\mathcal{P}_{i}$ are likely to occur after the collapse of the wave function.
@@ III Quantum entanglement production
$-\mathcal{P}_{k}=\mathcal{P}_{l}+\mathcal{P}_{m},$ (11)
where $k,l,m\in\{1,2,3\}$. Comparing these results with Eq. (5) for which
@@
$-\mathcal{P}_{j}(1+\beta\mathcal{P}_{j}^{2})=-P,$ (13)
@@
$|-P\rangle=\sum_{k=1}^{3}\gamma_{k}|-\mathcal{P}_{k}\rangle.$ (14)
@@
At this step, consider a two-particle system whose total momentum in the presence of minimal length (or equally, the total generalized momentum) is zero so that $P_{1}=-P_{2}\equiv P$ and they are moving along an arbitrary direction (such as the $x$ axis). Clearly, the momentum state $|P,-P\rangle$ indicates that the outcome of measuring $P_{i}$ is completely predictable. Interestingly enough, the generalized momentum conservation ($\sum_{j=1}^{j=2}P_{j}=0$) in the $\hat{P}-$space also inspires the canonical momentum conservation ($\sum_{j=1}^{j=2}p_{j}=$0) generating the $p_{1}=-p_{2}$ condition. Accordingly, if purely quantum mechanical momentum measurements are employed (or equally, the $|p\rangle$ space is focused on), then Eqs. (6) and (14) lead us to the state
$\sum_{i=1}^{3}\alpha_{i}\gamma_{i}|\mathcal{P}_{i},-\mathcal{P}_{i}\rangle,$ (15)
apart from a normalization constant. It unveils a novel quantum non-locality that exposes itself in the canonical momentum space. Therefore, the existence of minimal length seems capable to generate quantum non-locality.
@@ IV Summary and Concluding Remarks
The discovered degeneracy in $\langle x|P\rangle$ seems breakable using the canonical momentum operator. It also signals to us that minimal length (and thus GUP) may be an origin for the emergence of complex numbers in QM, which play impressive roles in various phenomena such as quantum non-locality AMJP ; Renou:2021dvp ; Li:2021uof ; Chen:2021ril ; Albert:2023kah ; Lancaster:2022vgi . Finally, and most importantly, calculations reveal the ability of minimal length to create quantum entanglement. The simplest spin versions of the EPR thought experiment Bohm:1957zz include a two-$1/2$ spin particle system whose spin state in the total spin space is either $|S=1,S_{z}=0\rangle$ or $|S=0,S_{z}=0\rangle$. The states’ decomposition using the spin state of each particle leads to the $\frac{1}{\sqrt{2}}\big[|+,-\rangle+|-,+\rangle\big]$ and $\frac{1}{\sqrt{2}}\big[|+,-\rangle-|-,+\rangle\big]$ states, respectively, addressing the quantum entanglement. The same behaviour is also obtained by expressing (measuring) $|P,-P\rangle$ in the canonical momentum bases (using canonical momentum operator) of each particle, leading to Eq. (15).
Acknowledgment
This work is based upon research funded by Iran National Science Foundation (INSF) under project No. 4046949.
Data availability
No data was used for the research described in the article.
Declaration of competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.